OpenStudy (anonymous):

Factor completely: 2x2 + 2x − 12

OpenStudy (anonymous):

Take out the common factor. Factor the remaining trinomial.\[2x^2 + 2x − 12=2(x^2+x-6)=2(x+3)(x-2)\]Do math every day.

OpenStudy (anonymous):

Thanks so much AnimalAin!

OpenStudy (anonymous):

No sweat. Do math every day.